Light bulbs with a stated mean lifetime of 750 hours have been stored in a warehouse for years . A sample of 10 of these bulbs is selected at random and destructively tested . The results of this sample gave a mean of 720 hours with a standard deviation of 40 hours . You wish to test whether prolong storage has significantly reduced the life expectancy of these light bulbs . You can assume that bulb lifetime is normally distributed. What is the correct p -value of this test ?
Answer #1

Given Mo = 750, ł 2720, 5= 40, 420105 Hoi u=750 Hai Me 750 720-750 401 o teat statistic = t= a Mo sho A ~ -2,372 df = 0-1=301